OpenStudy (anonymous):

I really need help with this please Let f(x)=2x+4x and g(x)=-6x-7. Find f(x)-g(x)

OpenStudy (anonymous):

@ash2326 can you help???

OpenStudy (jbo11):

The problem tells you what f(x) = which is (2x+4x) and g(x) is (-6x-7). Then subtract the two.

OpenStudy (anonymous):

wait thats all i have to do omg

OpenStudy (jbo11):

I would think so. So then its (2x+4x)-(-6x-7)=0

hartnn (hartnn):

just don't equate it to 0 (=0 )

OpenStudy (anonymous):

but wouldn't it be -7, because 2x+4x=6x and 6x-6x=0

hartnn (hartnn):

nopes, you are not subtracting correctly .... (2x+4x)-(-6x-7) distribute the - sign through the 2nd bracket 2x+4x - (-6x) - (-7) got this ?

OpenStudy (anonymous):

so it would be 12x+7?

hartnn (hartnn):

now you got it! :) correct.

OpenStudy (anonymous):

oh okay(: thank you(:

hartnn (hartnn):

welcome ^_^